find the equation for the line that contains (0,3)and is perpendicular to the line 6x-2y=1
Answer by mananth(16946)

6 x -2 y = 1
Find the slope of this line

-2 y = -6 x + 1
Divide by -2
y = 3 x + -0.5
Compare this equation with y=mx+b
slope m = 3.00

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be - 1/3

m= - 1/3 ,point ( 0 , 3 )
Find b by plugging the values of m & the point in
y=mx+b
3 = 0.00 + b
b= 3
m= - 1/3
The required equation is y = - 1/ 3 x + 3
m.ananth@hotmail.ca